# Lessons from pendulums: A design comparison of three lab activities

### v1

@@ II.1.2 Theoretically Informed
As one of the first evidence-based or research-based reformations of laboratory instruction, the RTP lab curriculum is a significant scholarly and historical milestone for PER as a field [may_historical_2023, otero_past_2017, thornton_tools_1987]. In addition to learning goals, Sokoloff et al. [sokoloff_realtime_2007] delineate several design principles “based on education research,” that undergird the RTP labs. The lab activities
  • are sequenced to provide students with a coherent observational basis for understanding a single topic area in one semester or quarter of laboratory sessions
  • provide activities that invite students to construct physical models based on observations and experiments
  • help students modify their common conceptions about physical phenomena that make it difficult for them to understand powerful general principles of physics
  • work well when performed in collaborative groups of two to four students
  • incorporate Microcomputer-based laboratory (MBL) tools so that students can test predictions by collecting and graphing data in real time
  • incorporate a learning cycle consisting of prediction, observation, comparison, analysis and quantitative experimentation
  • provide opportunities for class discussion of student ideas and findings and
  • integrate homework assignments designed to reinforce critical concepts and skills
The central innovation of the RTP labs are the MBL tools: sensors and interfaces that enable the real-time collection, display, and analysis of various physical quantities. These MBL tools were designed “with the student learner in mind” [thornton_tools_1987]. These tools enable students to focus on the physical world, giving students the ability to explore and learn from the world; the immediate feedback of the devices extends the range of potential investigations and the graphical display options can provide students a foothold in developing representational competence [thornton_learning_1990]. The design of the MBL tools as well as the RTP lab curriculum builds off of findings from cognitive science and education research at the time (e.g., [mcdermott_student_1987, halloun_common_1985, arons_achieving_1983]) as well as the 1998 AAPT laboratory recommendations [american_association_of_physics_teachers_goals_1998].
ISLE similarly aims for students both to learn concepts and to engage in empirical inquiry, but differs from RTP in how it implements experimentation. ISLE aims more explicitly at supporting students’ development in practices of empirical science. As part of that, it shifts responsibility for experimental design onto students, trading richer experience for the efficiency and reliability of successful, canonical findings. Like RTP, ISLE involves “cycles” of learning and experimentation, but the form differs. RTP emphasize student predictions and hypotheses, in part to elicit and respond to student misconceptions. ISLE, in contrast, begins with observational experiments and “something that needs explaining” [brookes_implementing_2020], in part for the designers’ views of student epistemologies as essential features of their engagement and targets for instruction.
To describe the theoretical foundations of their curriculum, Brookes et al. [brookes_implementing_2020] overview their “core intentionalities” [Garrison_MacMillan] – their beliefs and values about what their teaching is trying to accomplish – and use them as criteria for evaluating potential theoretical frameworks [Cobb2007]. With these criteria in mind, they discuss the “bricolage” of theoretical perspectives that actualize their beliefs about how learning happens and how to support it. This discussion expands on their 2010 paper [etkina_design_2010], which highlighted interpretive knowing as a desired form of cognition and activity that guided their design. Such theoretical descriptions not only provides specificity around important design decisions but also provides context for how and to what end the discussed theories were implemented.
In response to the need to teach remotely during the beginning of the COVID-19 pandemic, CU-Boulder redesigned their lab course to be a Course-based Undergraduate Research Experience (CURE) [werth_impacts_2022]. The authentic discovery that defines CUREs made it an attractive option for their redesign, as it supports engagement in authentic scientific practices, including collaboration and iteration; furthermore, the emphasis on authenticity (with respect to professional physics research) is expected to support motivation and persistence. In other words, authenticity as a theoretical construct shaped many of the design decisions in the construction of this curriculum.
The designers expect the authenticity of this research to support learning in two main ways: first, authentic research requires scientific practices like reading published literature, developing a research plan, data analysis, collaboration, and iteration; second, the authenticity of this research goes beyond “offer[ing] students glimpses of what it means to do experimental physics” [smith_best_2021], which can impact student motivation and affect [oliver_student_2023]. Oliver et al. [oliver_student_2023] identify six different aspects of the course designed to engage students in authentic research: a scaffolded literature review, meetings with the PI of the project (twice per semester for each group), metacognitive reflections, teamwork, peer review, a final reflection structured as a “memo to future researchers”, and the publication of their research.
The research from the first few semesters of this course highlights that engagement in scientific practices did occur [oliver_student_2023], students felt they participated in authentic research [werth_impacts_2022, oliver_student_2023], students felt that they had achieved their teamwork goals [werth_assessing_2022], and students felt that teamwork contributed to their success in the course [werth_assessing_2022].
@@
The theoretically-informed curricula described here all invoke learning theories, albeit in different ways: the design principles of RTP were inspired by educational research, although the exact connections are unclear; the researchers at CU Boulder outline expectations about how authenticity supports learning and list various features of their CURE expected to promote it; Brookes et al. [brookes_implementing_2020] list their overarching instructional values and intentionalities that help them select theoretical frameworks to do design work.
@@
As with learning goals, theory does not uniquely and completely determine design [Cobb2007, elby_rethinking_2020]. At the same time, these theoretically-informed approaches to description and explanation can provide insight into the design process and the links between learning goals and pedagogical structures. Explicating the reasoning undergirding design decisions, especially theoretically-grounded reasoning, can generate instructionally actionable insight [elby_rethinking_2020] and testable conjectures [disessa_cobb_2004, Cobb2007, sandoval2014].
@@ III Instructional Contexts
Although throughout this paper we refer to “our” theoretical commitments, we should be clear that these labs were designed and are implemented collaboratively. At Cornell, Emily Smith, Phil Krasicky, Mark Lory-Moran, Michael Niemack, Jared Maxson, Cristina Schlesier, and Rebeckah Fussell all contributed to the curriculum design, in addition to co-author Holmes. Holmes and her colleagues also were supported by an internal grant to do the research and development for Thinking Critically in Physics. [^3] https://www.physport.org/curricula/thinkingcritically/ At Tufts, Timothy Atherton and Hugh Gallagher contributed to the curriculum design, in addition to co-authors Hammer, Tobin, and Wagoner.
@@
The labs at both UWB and Cornell are stand-alone courses, and the lab curriculum at UWB is an adaptation of the Cornell curriculum. The Tufts lab discussed here is connected to the calculus-based introductory course taught by Hammer. Furthermore, we specifically discuss the spring instantiation of the Tufts labs; in Fall semesters the labs are shared across both the algebra-based and calculus-based courses, there are some changes to the lab curriculum discussed in this paper, and the overall course management proceeds differently.
The labs at Cornell are instructed by graduate TAs, they last two hours each, and there are a maximum of 20 students per section. At UWB, professors instruct the labs, which last two hours and have a maximum of 24 students. The Tufts labs are instructed by graduate and undergraduate TAs, last two and half hours, and there are a maximum of 14 students per section.
At both UWB and Tufts, the pendulum lab activity occurs over three lab sessions; at Cornell the activity takes three lab sessions. At Cornell and UWB, as stand-alone courses there is also a lecture portion of the lab course, where students reflect on their labs and work through activities to develop statistical skills. In contrast to Cornell and UWB, there is no assigned homework for students in the Tufts labs. Students turn in notes and a report, but time is allocated in lab for them to work on their report
@@ IV Three Versions of the Pendulum Labs
Our designs share foundations, orientations, and aspirations. The instructional labs each of us have had a hand in designing and implementing take as central a desire to engage students in doing of science. Specifically “what it means to do experimental physics: the approach, techniques, skills and ways of thinking when conducting authentic physics experiments” ([smith_best_2021], p. 1). Crucially, this doing must be active and intentional—it must be epistemically agentive. We design our labs in such a way to encourage and support students taking up a role as knowledge agents, who have the ability to creatively and intentionally pursue goals relevant to knowledge production in this classroom environment.
In the subsections that follow, we will provide a more in-depth description of the pedagogical logic. That is, an outline of the design itself alongside an explanation of the reasoning for the various design decisions. In these descriptions we seek to make clear how differences arise in our designs from different prioritizations and emphases of that overarching goal as well as different “secondary” considerations.
@@ IV.1 Cornell
At Cornell, the design of the pendulum creates a context for experimental practices to have purpose and relevance by setting students up to encounter an empirical-theoretical discrepancy. The instructional materials for the Cornell pendulum lab can be accessed through PhysPort. [^4] https://www.physport.org/curricula/thinkingcritically/ The lab instructions discuss the pendulum period equation derived from a model (with assumptions listed), and the guidance of the in-lab instructions set students on a path to produce data that indicates a deviation from this model. Although students are assessed on their process and the quality of their work, rather than by whether they arrive at the correct conclusion
At the heart of this lab activity is the epistemological richness of evaluating a model in conversation with experimental work; the (potential) conflict creates a need not only for calculating and interpreting uncertainty, but also for experimental design and data visualization skills that work together to respond to the theoretical-experimental discrepancy. This rich context aligns with the overarching course goal of students developing a more sophisticated understanding of what it means to know things and construct knowledge in physics.
As the first lab activity, the structure of this lab aims to cultivate student agency. The guidance of the in-lab instructions constrains the field of inquiry, but in ways that provide students with a sense of direction for their work so as to avoid the frustration of not knowing what to do. The instructions are open-ended to create decision-making opportunities. For example, in the first lab the experimentation begins with a direction to set up a pendulum and record period measurements at 10° and 20° but students are not told how carry out this experimentation. In this way, the in-lab instructions begin to shift responsibility for experimentation and knowledge production onto the students.
@@
After this discussion, there is one more thing students must do before they begin their experimentation: properly connect their CoLab or Jupyter Notebook to their utilities folder. The instructional team at Cornell recently updated their in-lab instructions to be written in Jupyter Notebook format, combining the in-lab instructions with students’ lab notebook. Students can use text cells to write out responses to questions or record other notes; the document also includes some base code that students can use to input and process their data. While coding and computational skills are not explicit foci of this course, including Jupyter Notebook/CoLab as a tool in this course allows students to become familiar with them and see them as useful scientific tools.
Once the code is properly set up, students begin their experimentation. Students start by constructing a pendulum and recording period measurements at 10° and 20°. They then perform some calculations (computing the mean, standard deviation, and standard uncertainty) and answer some questions connecting those statistical measures to the physical causes of variability and uncertainty in their data. This work launches them into the main focus of the first session: reducing uncertainty. Groups can choose between two options for their main experimental work of the session: exploring how increasing swings per trial reduces uncertainty or comparing repeated trials to multi-swing trials.
In between lab sessions, during the lecture portion of the course, students participate in a contrasting cases invention activity on quantifying distinguishability. The activity introduces students to the “t-prime” (t’) statistic [holmes_quantitative_2015], with the idea that the contrasting cases/invention structure supports student’s conceptual understanding of this statistical tool. In the following lab activity, students are prompted to use and interpret t’ values for the data they produced in the first session. The goal of the contrasting cases invention activity is to provide students with a tool they understand (as opposed to it being some black box) that they can use to make meaning from their data.
At the beginning of the second lab session, students compute t’ values for their data and are then asked about “reasonable next steps or new questions based on your set of t’ values?” The main focus of the second lab session is pushing models to their limits. As in the first session, groups can choose between two options, “reducing uncertainty” or “pushing on assumptions.” In both cases, students must explain how their task helps test the limits of the model and better understand the physical phenomenon. Groups generate ideas for possible experiments and then must “systematically test multiple measurement methods” for their chosen task.
@@
The directions and prompting to reduce uncertainty and test the limits of the model increase the likelihood that students encounter an empirical discrepancy. This discrepancy establishes a purpose for their experimental practices; in this context, ideally, students perceive a genuine need for careful consideration of uncertainty, statistical analyses, data visualization, and experimental design. At the same time, recognizing and responding to this discrepancy is not a requirement of the lab, and the assessment practices reflect that. The assessment emphasis is not on the content of their conclusions (whether they produce “the correct answer”) but rather on their process and explanation.
Instructors expect that this activity will surface and cue students’ expectations to verify authoritative knowledge and follow prescribed instructions. This particular approach aims to elicit and then respond to, if not confront, misaligned or improper expectations. At the end of the activity, through instructor feedback on their lab report and discussion in the associated lecture, students can reflect on their process and approach. Altogether, the lab activity aims to gradually transform students’ expectations about what instructional labs are about and cultivate their agency. During this lab unit and throughout the course, the specificity of the instructions is scaled back so that students take up more responsibility for figuring out what to do.
@@ IV.2 UWB
The lab curriculum at UWB is an adaptation of the Thinking Critically in Physics curriculum described above. The instructional materials for the UWD pendulum lab can be accessed through QUBES CourseSource [QUBES]. The design of this pendulum lab emphasizes meaningful, intentional use of experimental practices and the importance of collaboration and community feedback. The overarching goal, of this activity and the lab course in general, can be summarized as supporting students to engage in intentional, purposeful activity and perceive their work as consequential for producing knowledge about physical phenomena. Moreover, science is done with people, in community, with established (but transformable) social and epistemic norms. Collaboration, within groups and among the whole class, is a central feature of the curriculum.
@@
The experimentation begins with a single pendulum for the whole class, whose period students measure after it is released from 10° and 20°. Importantly, students are given no instructions on how to do so. Each individual measurement is recorded in a shared data table, and then the students (in their lab groups) are tasked with creating a histogram of these measurements to visualize the variability. This task also involves determining the precision of their measurement and calculating the mean and statistical uncertainty (standard deviation) of the data. This specific activity seeks to both direct student attention towards particular concepts (the role of data visualization and the physical meaning of uncertainty) and to foster student agency. Notably, the experimental question is neutral: instructors offer no theoretical context about simple harmonic motion or the underlying physics that would generate specific predictions.
@@
There is an agentive tradeoff occurring here: it is not the students’ choice to construct a histogram. At the same time, constructing a histogram involves choice and deliberation. Students have to figure out what an appropriate bin size is and what “appropriate” means in this context. They are responsible for making the measurements in the first place, and, once they construct a histogram, they are responsible for interpreting it. There is no predetermined correct answer their histogram is supposed to reveal; the purpose of the task is to do data interpretation and make sense of this phenomenon. Additionally, by directing students to create a histogram, they have a starting point or foothold in beginning to think about data visualization and the relationship between variation and uncertainty. The agentive tradeoff aims to forestall the inhibitive effects of discomfort or frustration with not knowing what to do and encourage choice and agency within the loose constraints of the activity.
@@
The second week begins with a short activity on comparing data sets quantitatively, which first asks lab groups to invent a way to distinguish example measurements and then summarizes some key points using real-life examples connected to the interests of the instructor (e.g., when Scherr teaches lab, she discusses datasets around classroom achievement in different instructional conditions). The invention activity primes students to notice and attend to particular quantities when interpreting and comparing data. Furthermore, as both an in-class, small group discussion activity and as part of their homework, students are asked “What are two data sets for which knowing whether they are the same or not would make a difference to your life or your community?” This reflection question aims to highlight the utility and meaningfulness of the data interpretation practices they will be engaging in during the second (and third) lab sessions.
@@
Following the mini-lecture and invention activity, the bulk of the experimental activity for the second lab revolves around reducing uncertainty. The students are tasked with using their data and experience to iteratively refine their experimental approach. The in-lab instructions direct students’ attention and activity towards certain concepts, such as measuring consecutive periods and collaborating with other groups, while also requiring explanations and interpretations of this work. Additionally, the goal of reducing uncertainty provides a way to show students that their activity is consequential and important in this learning environment: at the end of the first week, students are asked to brainstorm ideas about how to reduce experimental uncertainty in measuring pendulum periods and then in the second week they are told to pick one of those methods to implement. Many students do not seem to expect that their answers could inform their activity in the lab: they often list unhelpful options such as "eliminate human error" or "reduce air resistance." The follow up in the second week provides another opportunity to shift students’ expectations about how lab works: this is a place where their ideas are consequential.
@@
The third week begins with an introduction to the t’ statistic [holmes_quantitative_2015] that students then use to quantify the distinguishability of the data sets they produced in the previous week. Instructions prompt students to interpret their calculated t’ values and make decisions based on those interpretations: again, their activity is constrained, but students are still responsible for generating interpretations, decisions, and conclusions. A central aspect of this course is using tools, resources, or data in meaningful ways. By introducing the t’ statistic in this way and scaffolding it’s use during lab, students have multiple opportunities to use the t’ statistic to derive meaning from and interpret their data.
@@
Throughout, the lab instructions provide students with tools (e.g., a histogram template, a statistics vocabulary sheet, and the t’ statistic) and then ask them to make meaning through the use of those tools. Even if the students don’t initially recognize their work as meaningful, the repeated direction in the in-lab questions to improve their experiment reminds students that their decisions matter. In addition to these prompts, instructor interactions also emphasize the importance of intentionality. That is, the ways in which instructors engage, question, and respond to students reinforce the notion that experimental decisions and actions should have purpose and meaning behind them.
@@
Partway through the third lab session the instructor sets up a communal data table for groups to fill in with their period measurements, uncertainty values, t’ values, and conclusion. This provides a communal artifact that sparks inter-group discussion about results and conclusions.
After they’ve turned in their lab notes for grading, as part of their lab homework students are introduced to the simple harmonic model and associated period equation. The homework questions ask them to restate their conclusion about the distinguishability of periods at 10° and 20° and asks whether their conclusion aligns with the model; students are then asked to reflect on how knowing the “correct” result of an experiment can create bias, with a specific prompt to generate strategies to mitigate potential issues from this sort of bias.

### v2

@@ II.1.2 Theoretically Informed
The curricula described in this section all explicitly invoke learning theories (and related empirical research) as they describe their designs, albeit in different ways: the design principles of RTP were inspired by educational research; the researchers at CU Boulder outline expectations about how authenticity supports learning and list various features of their CURE expected to promote it; the designers of ISLE list their overarching instructional values and intentionalities that help them select theoretical frameworks to do design work.
As one of the first evidence-based or research-based reformations of laboratory instruction, the RTP lab curriculum is a significant scholarly and historical milestone for PER as a field [52, 55, 86]. In addition to learning goals, Sokoloff et al. [81] delineate design principles “based on education research,” that undergird the RTP labs. The lab activities
  • are sequenced to provide students with a coherent observational basis for understanding a single topic area in one semester or quarter of laboratory sessions;
  • provide activities that invite students to construct physical models based on observations and experiments;
  • help students modify their common conceptions about physical phenomena that make it difficult for them to understand powerful general principles of physics;
  • work well when performed in collaborative groups of two to four students;
  • incorporate Microcomputer-based laboratory (MBL) tools so that students can test predictions by collecting and graphing data in real time;
  • incorporate a learning cycle consisting of prediction, observation, comparison, analysis, and quantitative experimentation;
  • provide opportunities for class discussion of student ideas and findings; and
  • integrate homework assignments designed to reinforce critical concepts and skills.
The central innovation of the RTP labs are the MBL tools: sensors and interfaces that enable the real-time collection, display, and analysis of various physical measurements. These MBL tools were designed “with the student learner in mind” [86]. They enable students to focus on the physical world, giving them the ability to explore and learn from the world; the immediate feedback of the devices extends the range of potential investigations, and the graphical display options can provide students a start in developing representational competence [85]. The design of the MBL tools, along with the RTP lab curriculum, builds off of findings from cognitive science and education research at the time (e.g., [53, 27, 4]) as well as the 1998 AAPT laboratory recommendations [2].
ISLE similarly aims for students both to learn concepts and to engage in empirical inquiry, but differs from RTP in how it implements experimentation. ISLE aims more explicitly at supporting students’ development in practices of empirical science. As part of that, it shifts responsibility for experimental design onto students, trading richer experience for the efficiency and reliability of successful, canonical findings. Like RTP, ISLE involves “cycles” of learning and experimentation, but the form differs. RTP emphasizes student predictions and hypotheses, in part to elicit and respond to student misconceptions. ISLE, in contrast, begins with observational experiments and “something that needs explaining” [9], in part for the designers’ views of student epistemologies as essential features of their engagement and targets for instruction.
To describe the theoretical foundations of their curriculum, Brookes et al. [9] overview their “core intentionalities” [51] – their beliefs and values about what their teaching is trying to accomplish – and use them as criteria for evaluating potential theoretical frameworks [10]. With these criteria in mind, they discuss the “bricolage” of theoretical perspectives that actualize their beliefs about how learning happens and how to support it. This discussion expands on their 2010 paper [19], which highlighted interpretive knowing as a desired form of cognition and activity that guided their design. Such theoretical descriptions provide not only specificity around important design decisions but also context for how and to what end the theories were implemented.
In response to the need to teach remotely during the beginning of the COVID-19 pandemic, CU-Boulder redesigned their lab course to be a Course-based Undergraduate Research Experience (CURE) [91]. The authentic discovery that defines CUREs made it an attractive option for their redesign, as it supports engagement in key scientific practices, including collaboration and iteration; furthermore, the emphasis on authenticity (with respect to professional physics research) is expected to support motivation and persistence. In other words, authenticity as a theoretical construct shaped many of the design decisions in the construction of this curriculum.
The designers expected the authenticity of this research to support learning in two main ways: first, authentic research requires scientific practices such as reading published literature, developing a research plan, data analysis, collaboration, and iteration; second, the authenticity of this research goes beyond “offer[ing] students glimpses of what it means to do experimental physics” [78], which is expected to positively impact student motivation and affect [54]. Oliver et al. [54] identify six different aspects of the course designed to engage students in such authentic research: a scaffolded literature review, meetings with the PI of the project (twice per semester for each group), metacognitive reflections, teamwork, peer review, a final reflection structured as a “memo to future researchers,” and the publication of their research.
The education research from the first few semesters of this course highlights that engagement in scientific practices did occur [54]. Students felt they participated in authentic research [91, 54], that they had achieved their teamwork goals [90], and that teamwork contributed to their success in the course [90].
@@
As with learning goals, theory does not uniquely and completely determine design [10, 18, 7]. Still, theoretically-grounded descriptions provide insight into the design process and the reasoning that links learning goals and pedagogical structures. Empirical study is necessary to understand what works when, for whom, and how, but the application of those findings to new situations and curricula necessarily involves theorizing. Explicating the reasoning for design decisions, especially theoretically-grounded reasoning, can generate instructionally actionable insight [18, 65] and empirically testable conjectures [16, 10, 69].
@@ III Instructional Contexts
Although throughout this paper we refer to the authors’ theoretical commitments, we should be clear that these labs were designed and are implemented collaboratively. At Cornell, Emily Smith, Phil Krasicky, Mark Lory-Moran, Michael Niemack, Jared Maxson, Cristina Schlesier, and Rebeckah Fussell all contributed to the curriculum design, in addition to co-author Holmes. Holmes and her colleagues were supported by an internal grant to do the research and development for Thinking Critically in Physics. [^3] https://www.physport.org/curricula/thinkingcritically/ At Tufts, Timothy Atherton and Hugh Gallagher contributed to the curriculum design, in addition to co-authors Hammer, Tobin, and Wagoner.
@@
The introductory lab at Cornell is a stand-alone course, and the lab curriculum at UWB is an adaptation of the Cornell curriculum. Each includes a lecture, where students reflect on their labs and work through activities to develop statistical skills. At Tufts and UWB, introductory labs are sections connected to the main lecture courses. Tufts has no lab-specific lecture and lab instructions change in small ways from one semester to the next. Here we focus on the spring instantiation of the Tufts labs, taught by Hammer, specifically from the spring 2024. In Fall semesters, the labs are shared across algebra-based and calculus-based courses, and the overall course management proceeds differently.
The labs at Cornell are instructed by graduate TAs, they last two hours each, and there are a maximum of 20 students per section. At UWB, professors instruct the labs, which last two hours and have a maximum of 24 students per section. The Tufts labs are instructed by graduate and undergraduate TAs, last two and half hours, and there are a maximum of 14 students per section.
At both UWB and Tufts, the pendulum lab activity occurs over three lab sessions; at Cornell the activity takes two lab sessions. The lab courses at Cornell and UWB assign homework; Tufts labs do not (students complete all notes and reports within the 2.5-hour sessions).
@@ IV Three Versions of the Pendulum Labs
Our designs share foundations, orientations, and aspirations. All take as central a desire to engage students in doing empirical science, focusing on “what it means to do experimental physics: the approach, techniques, skills and ways of thinking when conducting authentic physics experiments” ([78], p. 1). Crucially, this doing must be active and intentional—it must be epistemically agentive. We design our labs in such a way as to encourage and support students taking up roles as knowledge agents, who have the ability to creatively and intentionally pursue goals relevant to knowledge production.
In the subsections that follow, we provide a more in-depth description of the pedagogical logic, an outline of the design itself along with an explanation of the reasoning for the various design decisions. In these descriptions, we seek to make clear how differences arise in our designs from different prioritizations and emphases of that overarching goal as well as different “secondary” considerations.
@@ IV.1 Cornell
At Cornell, the design of the pendulum activity creates a context for experimental practices to have purpose and relevance by setting students up to encounter an empirical-theoretical discrepancy. The instructional materials for the Cornell pendulum lab can be accessed through PhysPort. [^4] https://www.physport.org/curricula/thinkingcritically/ The lab instructions discuss the pendulum period equation derived from a model (with assumptions listed), and the guidance of the in-lab instructions set students on a path to produce data that indicates a deviation from this model. Students are assessed on their process and the quality of their work, rather than by whether they arrive at the correct conclusion.
At the heart of this lab activity is the epistemological richness of evaluating a model in conversation with experimental work; the (potential) conflict creates a need not only for calculating and interpreting uncertainty, but also for experimental design and data visualization skills that work together to respond to the theoretical-experimental discrepancy. This rich context aligns with the overarching course goal of students developing a more sophisticated understanding of what it means to know things and to construct knowledge in physics.
As the first lab activity, the structure of this lab aims to cultivate student agency. The guidance of the in-lab instructions constrains the field of inquiry, but in ways that provide students with a sense of direction for their work so as to avoid the frustration of not knowing what to do. The instructions are open-ended to create decision-making opportunities. For example, in the first lab the experimentation begins with a direction to set up a pendulum and record period measurements at 10° and 20° but students are not told how to carry out this experimentation. In this way, the in-lab instructions begin to shift responsibility for experimentation and knowledge production onto the students.
@@
After this discussion, there is one more thing students must do before they begin their experimentation: properly connect their CoLab or Jupyter Notebook to their utilities folder. The instructional team at Cornell recently updated their in-lab instructions to be written in Jupyter Notebook format, combining the in-lab instructions with students’ lab notebook. Students can use text cells to write out responses to questions or record other notes; the document also includes some base code that students can use to input and process their data. While coding and computational skills are not explicit foci of this course, including Jupyter Notebook/CoLab as tools in this course allows students to become familiar with them and see them as useful.
Once the code is properly set up, students begin their experimentation. Students start by constructing a pendulum and recording period measurements at 10° and 20°. They then perform some calculations (computing the mean, standard deviation, and standard uncertainty) and answer some questions connecting those statistical measures to the physical causes of variability and uncertainty in their data. This work launches them into the main focus of the first session: reducing uncertainty. Groups can choose between two options for their main experimental work of the session: exploring how increasing swings per trial reduces uncertainty or comparing the approaches of using repeated single swing trials to a multi-swing trial.
In between lab sessions, during the lecture portion of the course, students participate in a contrasting cases invention activity on quantifying distinguishability. The activity introduces students to the “t-prime” (t’) statistic [31], similar to Student’s t-test, with the idea that the contrasting cases/invention structure supports student’s conceptual understanding of this statistical tool. In the following lab activity, students are prompted to use and interpret t’ values for the data they produced in the first session. The goal of the contrasting cases invention activity is to provide students with a tool they understand (as opposed to a black box) that they can use to make meaning from their data.
At the beginning of the second lab session, students compute t’ values for their data and are then asked about “reasonable next steps or new questions based on your set of t’ values?” The main focus of the second lab session is pushing models to their limits, and they are expected to use t’ to help them in that work. As in the first session, groups can choose between two options; the two options this week are “reducing uncertainty” or “pushing on assumptions.” In both cases, students must explain how their task helps test the limits of the model and better understand the physical phenomenon. All groups generate ideas for possible experiments and then must “systematically test multiple measurement methods” for their chosen task.
@@
The directions and prompts to reduce uncertainty and test the limits of the model increase the likelihood that students encounter an empirical discrepancy. This discrepancy establishes a purpose for their experimental practices; in this context, ideally, students perceive a genuine need for careful consideration of uncertainty, statistical analysis, data visualization, and experimental design. At the same time, recognizing and responding to this discrepancy is not a requirement of the lab, and the assessment practices reflect that. The assessment emphasis is not on the content of their conclusions (whether they produce “the correct answer”) but rather on their process and explanation.
Instructors expect that this activity will surface and cue students’ expectations to verify authoritative knowledge and follow prescribed instructions. This approach aims to elicit and then respond to, if not confront, misaligned or improper expectations. At the end of the activity, through instructor feedback on their lab report and discussion in the associated lecture, students can reflect on their process and approach. Altogether, the lab activity aims to gradually transform students’ expectations about what instructional labs are about and cultivate their agency. During this lab unit and throughout the course, the specificity of the instructions is scaled back so that students take up more responsibility for figuring out what to do.
@@ IV.2 UWB
The lab curriculum at UWB is an adaptation of the Thinking Critically in Physics curriculum described above. The instructional materials for the UWB pendulum lab can be accessed through QUBES CourseSource [72]. The design of this pendulum lab emphasizes meaningful, intentional use of experimental practices and the importance of collaboration and community feedback. The overarching goal, of this activity and the lab course in general, can be summarized as supporting students to engage in intentional, purposeful activity and perceive their work as consequential for producing knowledge about physical phenomena. Moreover, science is done with people, in community, with established (but transformable) social and epistemic norms. Collaboration, within groups and among the whole class, is a central feature of the curriculum.
@@
The experimentation begins with a single pendulum for the whole class, whose period students measure after it is released from 10° and 20°. Importantly, students are given no instructions on how to do so. They are also are asked to make a prediction about how the periods at 10° and 20° will compare, "based on the physics that you know." Each individual measurement is recorded in a shared data table, and then the students (in their lab groups) are tasked with creating a histogram of these measurements to visualize the variability. This task also involves determining the precision of their measurement and calculating the mean and statistical uncertainty (standard deviation) of the data. This specific activity seeks to both direct student attention towards particular concepts (the role of data visualization and the physical meaning of uncertainty) and to foster student agency.
@@
There is a tradeoff pertaining to students’ agency here: it is not the students’ choice to construct a histogram. At the same time, constructing a histogram involves choice and deliberation. Students have to figure out what an appropriate bin size is and what “appropriate” means in this context. They are responsible for making the measurements in the first place, and, once they construct a histogram, they are responsible for interpreting it. There is no predetermined correct answer their histogram is supposed to reveal; the purpose of the task is to do data interpretation and make sense of this phenomenon. Additionally, by directing students to create a histogram, they have a starting point or foothold in beginning to think about data visualization and the relationship between variation and uncertainty. The agentive tradeoff aims to forestall the inhibitive effects of discomfort or frustration with not knowing what to do and encourage choice and agency within the loose constraints of the activity.
@@
The second week begins with a short activity on comparing data sets quantitatively, which first asks lab groups to invent a way to distinguish example measurements and then summarizes some key points using real-life examples connected to the interests of the instructor (e.g., when Scherr teaches lab, she discusses datasets around classroom achievement in different instructional conditions). The invention activity primes students to notice and attend to particular quantities when interpreting and comparing data. Furthermore, both as an in-class, small group discussion and as part of their homework, students are asked “What are two data sets for which knowing whether they are the same or not would make a difference to your life or your community?” This reflection question aims to highlight the utility and meaningfulness of the data interpretation practices they will be engaging in during the second (and third) lab sessions.
@@
Following the mini-lecture and invention activity, the bulk of the experimental activity for the second lab revolves around reducing uncertainty. The students are tasked with using their data and experience to iteratively refine their experimental approach. The in-lab instructions direct students’ attention and activity towards certain concepts, such as measuring consecutive periods and collaborating with other groups, while also requiring explanations and interpretations of this work.
Additionally, the goal of reducing uncertainty provides a way to show students that their activity is consequential and important in this learning environment: at the end of the first week, students are asked to brainstorm ideas about how to reduce experimental uncertainty in measuring pendulum periods and then in the second week they are told to pick one of those methods to implement. Many students do not seem to expect that their answers could inform their activity in the lab: they often list unhelpful options such as "eliminate human error" or "reduce air resistance." The follow up in the second week provides another opportunity for their coming to see lab as a place where their ideas are consequential.
@@
The third week begins with an introduction to the t’ statistic [31], which the students apply to the data sets they produced in the previous week. Instructions prompt students to interpret their calculated t’ values and make decisions based on those interpretations. This again constrains their activity, but they remain responsible for generating interpretations, decisions, and conclusions. By introducing the t’ statistic in this way and scaffolding its use, the lab activities provide students multiple opportunities to use the t’ statistic to derive meaning from and interpret their data.
@@
Throughout, the lab instructions provide students with tools (e.g., a histogram template, a statistics vocabulary sheet, and the t’ statistic) and then ask them to make meaning through their use. Even if the students do not initially recognize these tools or their use of them as meaningful, the repeated written prompts direction to improve their experiment work to shift them in that direction. During lab meetings, the ways in which instructors engage, question, and respond to students similarly work to reinforce this framing, that the students’ use of these tool should have purpose and meaning.
@@
The experimental prompt is neutral with respect to outcome: It provides no theoretical context about simple harmonic motion or Galileo’s claims that favor specific predictions. Partway through the third lab session, the instructor sets up a shared data table for groups to fill in with their period measurements, uncertainty values, t’ values, and conclusion. This provides a communal artifact that sparks inter-group discussion about results and conclusions.
It is not until after they have turned in their lab notes for grading that students are introduced to the simple harmonic model and associated period equation. Homework questions ask them, first, to restate their conclusion about the distinguishability of periods at 10° and 20°, and second, whether their conclusion aligns with the model. They are then asked to reflect on how knowing the “correct” result of an experiment can create bias, with a specific prompt to generate strategies to mitigate it.